ISRAEL JOURNAL OF MATHEMATICS

TABLE OF CONTENTS, VOLUME 208, 2015

Adiprasito, K. A. and I. Izmestiev Derived subdivisions make every PL sphere polytopal

443

Aldecoa, R. T. d. The absolute continuous spectrum of skew products of compact Lie groups

323

Altınel, T., J. Burdges and O. Frécon Structure of Borel subgroups in simple groups of finite Morley rank

101

Bagaria, J., M. Magidor and H. Sakai Reflection and indescribability in the constructible universe

1

Berarducci, A. and M. Mamino Groups definable in two orthogonal sorts

413

Berele, A. and S. Catoiu Asymptotic behavior of S_{n} degrees associated with rational functions

351

Berkovich, Y. On the theorem of Mann about p-groups all of whose nonnormal subgroups are elementary abelian

451

Brieussel, J. and R. Tanaka Discrete random walks on the group Sol

291

Burdges, J. See T. Altınel Catoiu, S. See A. Berele Chen, Q., Y. Wu and B. Xu On one-dimensional and singular Cal-

abi’s extremal metrics whose Gauss curvatures have nonzero umbilical Hessians

385

Dayan, Y. and E. Glasner Hypercyclic operators, Gauss measures and Polish dynamical systems

79

Fernández-Alcober, G. A., L. Legarreta, A. Tortora and M. Tota Some restrictions on normalizers or centralizers in finite pgroups

193

Fernández-Polo, F. J., A. M. Peralta and M. I. Ramírez A Kadec–Pelczyński dichotomy-type theorem for preduals of JBW^{∗}- algebras

45

Frécon, O. See T. Altınel Glasner, E. See Y. Dayan Izmestiev, I. See K. A. Adiprasito Kharchenko, V. K. Explicit coproduct formula for quantum groups

of infinite series

13

Kuhlmann, F. See K. Kuhlmann Kuhlmann, K., F. Kuhlmann and S. Shelah Symmetrically com-

plete ordered sets abelian groups and fields

261

López-Permouth, S. R., J. Moore, N. Pilewski and S. Szabo Algebras having bases that consist solely of units

461

Legarreta, L. See G. A. Fernández-Alcober Márki, L., J. Meyer, J. Szigeti and L. v. Wyk Matrix repre-

sentations of finitely generated Grassmann algebras and some consequences

373

Magidor, M. See J. Bagaria Mamino, M. See A. Berarducci Meyer, J. See L. Márki

Meyer, M., C. Schütt and E. M. Werner Affine invariant points  163

Moore, J. See S. R. López-Permouth Panagiotopoulos, A. Extendability of automorphisms of generic

substructures

483

Pazzis, C. d. S. The classification of large spaces of matrices with bounded rank

219

Peralta, A. M. See F. J. Fernández-Polo Pilewski, N. See S. R. López-Permouth Ramírez, M. I. See F. J. Fernández-Polo Sakai, H. See J. Bagaria Schütt, C. See M. Meyer Shelah, S. See K. Kuhlmann Szabo, S. See S. R. López-Permouth Szigeti, J. See L. Márki Tanaka, R. See J. Brieussel Tortora, A. See G. A. Fernández-Alcober Tota, M. See G. A. Fernández-Alcober Werner, E. M. See M. Meyer Wu, Y. See Q. Chen Wyk, L. v. See L. Márki Xu, B. See Q. Chen